OpenStudy (anonymous):

Where 0<=x<360, Solve the equation

OpenStudy (anonymous):

\[3\sin x-5+\csc x=0\]

OpenStudy (mertsj):

\[3\sin ^2x-5\sin x+1=0\]

OpenStudy (anonymous):

But I simplified that to \[\sin ^{2}x-3\sin x+1\]

OpenStudy (mertsj):

I don't see how you get that?

OpenStudy (anonymous):

Give me a sec

OpenStudy (anonymous):

Oh wait.. I was looking at something else. My bad.

OpenStudy (anonymous):

Would I use the quadratic formula to find the roots of 3sin^2x−5sinx+1=0?

OpenStudy (mertsj):

\[3sinx \times sinx-5\times sinx+\frac{1}{sinx}\times sinx=0 \times sinx\]

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

\[\frac{ 3\pm \sqrt{5}}{ 2 }\] But then what? I can't isolate for x, can I?

OpenStudy (mertsj):

sin x = (3+sqrt5)/2 or sin x = (3-sqrt5)/2

OpenStudy (anonymous):

So would I find the arcsin of (3+sqrt5)/2 and (3-sqrt5)/2 for x?

OpenStudy (mertsj):

Can't be the first one because sin can't be bigger than 1.

OpenStudy (mertsj):

so sin x = .3820

OpenStudy (anonymous):

Does there happen to be a way I could solve this without using a calculator?

OpenStudy (mertsj):

If you have a table of values you could.

OpenStudy (mertsj):

Did you get 22.5 degrees and 157.5 degrees?

OpenStudy (anonymous):

Do we need my assistance here or are we ok?